The equationF=d/dt(m(t)v(t)) as applied to a variable mass system is true when Choose at least one answer. 1. mass enters the variable mass system with zero velocity relative to a stationary observer 2. mass leaves the system with zero velocity relative to a stationary observer 3. mass enters the system with zero velocity relative to the variable mass system 4.mass leaves the system with zero velocity relative to the variable mass system

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

F=d/dt(m(t)v(t))=m(t)dv(t)/dt +v(t)dm(t)/dt, therefore (1) and (2) are correct

'Win exciting gifts by answering the questions on Discussion Forum.'